OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

hartnn (hartnn):

in the equation y=mx+c m =slope = 4 y= 4x+c plug x = -3, y = -1 in this to find the value of c

OpenStudy (anonymous):

thank you!

hartnn (hartnn):

welcome ^_^ if you want me to verify, tell me what u get for 'c' :)

OpenStudy (anonymous):

tell me what you got and ill tell you what i got

hartnn (hartnn):

tell me what you got and ill tell you what i got

OpenStudy (anonymous):

thats lame man

OpenStudy (anonymous):

idk how to do this ok im only 10

OpenStudy (anonymous):

plz help meh

hartnn (hartnn):

y= 4x+c -1 = 4(-3) +c what is 4 (-3) =.. ?

OpenStudy (anonymous):

pliiiiz help

hartnn (hartnn):

relax, i am helping :)

OpenStudy (anonymous):

hold on while i pull out my calculator

OpenStudy (anonymous):

-12

hartnn (hartnn):

yes, so, -1 = -12 + c add 12 on both sides of the equation, what do u get ?

OpenStudy (anonymous):

-13=12+c

hartnn (hartnn):

how ?

hartnn (hartnn):

-1 + 12 = -12 + 12 +c

OpenStudy (anonymous):

idk gosh im onlya youngen

hartnn (hartnn):

whats -12 + 12 ? -1 + 12 ?

OpenStudy (anonymous):

0 and then 11

hartnn (hartnn):

yes so, 11 = 0+c so, c= 11 got this ?

OpenStudy (anonymous):

ya i feel you

OpenStudy (anonymous):

but that isnt an answer choice

hartnn (hartnn):

it is c= 11 y = 4x+c gives y= 4x+11

OpenStudy (anonymous):

thats what i was about to say lol thanks for your help dude

hartnn (hartnn):

welcome ^_^